# Threshold-less and Flexibly Tunable Frequency Comb via Floquet Engineering

### v2

@@ Results
To achieve a stable self-induced oscillation, we applied a fixed drive power $P_{\mathrm{d}}=10$ dBm. As shown in Fig. 3a, we send a pump tone to cavity-2 and characterize the emergence and evolution of the frequency comb in the cavity response signal with different detuning $\Delta_{\mathrm{p}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c2}}$ and power $P_{\mathrm{p}}$ (see Fig. 3b). The local minima occurring at the sideband frequencies $\Delta_{\mathrm{p}}=m\omega_{\mathrm{m}}$ are due to the fact that the pump frequency matches one of the Floquet sidebands, thereby allowing efficient photon injection. We demonstrate the flexible tuning capability of the Floquet frequency comb within the range of $-3\omega_{\mathrm{m}}\leq\Delta_{\mathrm{p}}\leq 3\omega_{\mathrm{m}}$. Notably, frequency comb can also emerge in the red sideband region of the cavity-2 ($\Delta_{\mathrm{p}}<0$), which is completely absent in Kerr optomechanical combs (for $\kappa<\Omega$) [47, 48, 49, 50, 57]. In theory, the pump power could be much lower, but it is constrained by the limitations of the detection accuracy in our experimental setup. Fig. 3c shows an example of the frequency comb at $\Delta_{\mathrm{p}}=-1\times\omega_{\mathrm{m}}$. As the pump power increases, the number of detectable frequency comb teeth increases, but the pump power itself does not affect the emergence of the frequency comb.
@@
For comparison, we also independently characterized the emergence and evolution of the Kerr optomechanical frequency comb in cavity-2. As shown by the red data points in Fig. 3b, the minimum threshold power $P_{\mathrm{p-Kerr}}=11\ \mathrm{dBm}$ and occurs at the blue sideband $\Delta_{\mathrm{p}}\approx\Omega_{\mathrm{m}}$ because of maximized antidamping, triggering the optomechanical instability most efficiently.
@@
Focusing on the Floquet mode of $m=1$ in Fig. 3b, the minimum generated power for the Floquet frequency comb is approximately reduced by six orders of magnitude compared to Kerr optomechanical combs. Moreover, for the $m=1$ mode, the width of the local minimum is significantly broader and shifts toward $m=0$. This occurs because, at this point, the pump not only serves as the probe of Floquet frequency comb, but also drives the mechanical oscillator acting similarly to the mechanism in Kerr optomechanical combs. The interaction between the pump signal and the mechanical oscillator results in a broader local minimum and the shift of this mode.
Effect of modulation strength
@@
Figure 4: a Effect of drive and pump power on Floquet frequency comb generation at $m=-1,0,+1$(Red, purple, blue). The colored regions indicate the Floquet comb is generated, whereas the white areas represent the absence of a detectable frequency comb. The green dashed line represents the power range selected for subsequent analysis. b Example frequency comb for $m=-1$: frequency domain, time domain signals (Red: real part; Blue: imaginary part) and phase space trajectories. c Effect of modulation strength on the relative magnitudes of the frequency comb teeth, with error bars representing standard deviations. The gradient-colored points represent the peak values of the corresponding $m$-th comb tooth.
@@
We investigate the impact of modulation strength on the Floquet frequency comb by setting the pump tone at $\Delta_{\mathrm{p}}/\omega_{\mathrm{m}}=-1,0,+1$ to match the $m$-th Floquet mode. As shown in Fig. 4a, while varying both the drive power and pump power, for $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$, the frequency comb is generated only when the Floquet modulation is activated, $P_{\mathrm{d}}>P_{\mathrm{dc}}$. The fluctuations in the minimum power value are attributed to measurement uncertainties. Notably, for $\Delta_{\mathrm{p}}=0$, when the pump power is sufficiently high, the frequency comb is generated before $P_{\mathrm{dc}}$. This is the result of the cooperative action of the optomechanical frequency comb under the external driving of two cavities. The results for $\Delta_{\mathrm{p}}=\omega_{\mathrm{m}}$ exhibit similar behavior. As at sufficiently high pump power, cavity-2 can independently generate a frequency comb via Kerr optomechanical interactions. Thus, we observe a gradual onset rather than a sharp transition with $\Delta_{\mathrm{p}}=\omega_{\mathrm{m}}$.
To minimize interference from the Kerr optomechanical frequency comb in our analysis, we focus on the pump at $\Delta_{\mathrm{p}}=-\omega_{m}$. Fig. 4b shows an example of Floquet frequency comb with $P_{\mathrm{d}}=11\,\mathrm{dBm}$ and $P_{\mathrm{p}}=0\,\mathrm{dBm}$. Time-domain signals and phase space trajectories indicate that the output comb maintains excellent coherence. The periodicity of the time-domain signal results from a 2 MHz offset between the demodulation frequency and the strongest peak, demonstrating the stability and coherence of the Floquet frequency comb in this regime. For the $\Delta_{\mathrm{p}}=-\omega_{m}$ detection, the resulting frequency comb intensity is not perfectly symmetric, which is attributed to the non-zero $\Delta$ term and $X_{\mathrm{0}}$ (see Supplementary Note V).
@@
We further demonstrate the effect of modulation strength for the $\Delta_{\mathrm{p}}=-\omega_{m}$ case with $P_{\mathrm{p}}=0\,\mathrm{dBm}$, shown in Fig. 4c. As the modulation strength increases, the Floquet cavity’s ability to scatter photons into higher-order sidebands strengthens, resulting in a decrease in the intensity of the peak at $m=0$ and an increase in the intensities of peaks at $m=\pm 1,\pm 2,\dots$. With further increase in modulation strength, the intensity of the lower-order sidebands stabilizes and begins to decrease, while the effect of higher-order sidebands becomes more pronounced. At higher $P_{\mathrm{d}}$ values, it is noteworthy that the intensity of $m=-1$ exhibits an anomalous increase. This deviation is due to the $m=-1$ mode overlapping with the pump signal, preventing it from following the expected intensity decrease in sync with $m=+1$. Further details of the frequency combs and numerical simulations can be found in Supplementary Note V.
@@ Discussion
We have proposed and demonstrated a new type of frequency comb via Floquet engineering. The approach developed in this work not only offers a threshold-less route for frequency comb emission but also significantly broadens the frequency tuning range. This advancement paves the way for the efficient generation and control of frequency combs, making them more suitable for low-power regimes required by sensitive systems. For example, in superconducting quantum computing devices, components such as superconducting resonators and qubits cannot operate under high photon inputs and cryogenic amplifiers have typically low saturation powers. Moreover, the generation of Floquet frequency comb is not limited to microwave optomechanical systems. Due to the extent to which the mechanical oscillator is driven, our demonstration shows up to approximately ten characteristic comb teeth. By optimizing stronger modulation sources, the number of frequency comb teeth can be rapidly expanded, with potential for self-referencing. Floquet frequency combs can be realized and applied in various systems exhibiting periodic modulation, such as SQUIDs [58], trapped ions [29], time crystals [59], and molecular or atomic systems [60], particularly in the optical domain. This holds promise for extending the capabilities of frequency combs in high dimensional quantum states [61], entanglement generation [36, 37], and integrated multi-frequency signal sources across more systems [58, 29].
@@ Methods
here $\psi_{n}(t)=\langle n|\psi,t\rangle$. Combined with Eq. (3), the steady-state solution can be expressed as
@@
$\left|u_{n},t\right\rangle$ $=e^{-in\Omega A\sin\Omega t}|n\rangle.$
We define $\left|u_{nm},t\right\rangle=\left|u_{n},t\right\rangle e^{im\Omega t}=u_{nm}(t)|n\rangle$, with $u_{nm}(t)=e^{-in\Omega A\sin\Omega t}e^{im\Omega t}$, which satisfies
@@
At this point, we have identified Brillouin regions over the frequency domain, where the single Brillouin region have a width of $\Omega$, and the different Brillouin regions are represented by the index $m$, called the Floquet index. And where $H_{F}=H(t)-i\partial_{t}$ is called the Floquet Hamiltonian, $\left|u_{nm},t\right\rangle$ is the time-periodic Floquet state, and quasi-energy $\varepsilon_{nm}=n\omega_{0}+m\Omega$. The $\left|u_{n},t\right\rangle$ can also be written as $\left|u_{n0},t\right\rangle$. Under the complete basis vector set $\left.\left\{\left|n_{mn}\right\rangle\right\rangle\right\}$ with $\left|n_{m}\right\rangle=|n\rangle e^{im\Omega t}$, the matrix elements of the Floquet Hamiltonian are
@@
where $H^{[0]}=\sum_{n}n\omega_{c}|n\rangle\langle n|$, $H^{[\pm 1]}=\sum_{n}n\frac{A}{2}|n\rangle\langle n|$, and $H^{[l]}=0$ for $|l|>1$. Thus, the Eq. (7) can be further expressed as
@@
$H_{F}^{(n)}=\left[\begin{array}[]{ccccc}\ddots&\vdots&\vdots&\vdots&\begin{sideways}$\ddots$\end{sideways}\\ \cdots&n\omega_{0}-\Omega&n\frac{A}{2}&0&\cdots\\ \cdots&n\frac{A}{2}&n\omega_{0}&n\frac{A}{2}&\cdots\\ \cdots&0&n\frac{A}{2}&n\omega_{0}+\Omega&\cdots\\ \begin{sideways}$\ddots$\end{sideways}&\vdots&\vdots&\vdots&\ddots\end{array}\right]\text{, }$ (10)
which satisfies $H_{F}^{(n)}\left|u_{nm}\right\rangle\rangle=\varepsilon_{nm}\left|u_{nm}\right\rangle\rangle$ according to the Eq. (6). It can be observed that each Fock state energy level splits into $12m+$ 1 sublevels, consistent with the periodic nature of the Floquet Hamiltonian. The off-diagonal elements, proportional to $A/2$, characterize the coupling strength between adjacent Floquet states, thereby determining the transition rates between different energy levels. This coupling reflects the influence of the periodic driving field and results in transitions mediated by the exchange of discrete energy quanta of size $\Omega$. For more detailed derivations, please refer to the Supplementary Note I.B.
@@
In the blue-detuned regime, the optomechanical damping rate $\gamma_{\mathrm{opt}}$ is negative. Initially, this results in an increase in the oscillator’s effective temperature. As the overall damping rate $\gamma_{\mathrm{m}}+\gamma_{\mathrm{opt}}$ becomes negative, any small initial fluctuation will grow exponentially over time, eventually reaching a steady-state regime. These are referred to as self-induced optomechanical oscillations (or phonon lasing regime), where the mechanical oscillation amplitude stabilizes at a fixed value $X_{\mathrm{m}}$. The governing dynamical equations are given by
@@
Ignoring in the initial phase $\phi_{\mathrm{m}}$, we have $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t)+X_{\mathrm{0}}$. In steady state, we define an amplitude-dependent effective optomechanical damping rate as
@@
[…]
@@
and the global phase $\varphi(t)=(g_{\mathrm{01}}X_{\mathrm{m}}/x_{\mathrm{zpf}}\omega_{\mathrm{m}})\sin(\omega_{\mathrm{m}}t)$. By substituting this Fourier series into Eq. (14), we obtain the implicit equation for $X_{\mathrm{0}}$ and $X_{\mathrm{m}}$:
[…]
@@
All other data relevant to this study are available from the corresponding authors. Source data are provided with this paper.

### v3

@@ Results
To achieve a stable self-induced oscillation, we applied a fixed drive power $P_{\mathrm{d}}=10$ dBm. As shown in Fig. 3a, we send a pump tone to cavity-2 and characterize the emergence and evolution of the frequency comb in the cavity response signal with different detuning $\Delta_{\mathrm{p}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c2}}$ and power $P_{\mathrm{p}}$ (see Fig. 3b). The local minima occurring at the sideband frequencies $\Delta_{\mathrm{p}}=m\omega_{\mathrm{m}}$ are since the pump frequency matches one of the Floquet sidebands, thereby allowing efficient photon injection. We demonstrate the flexible tuning capability of the Floquet frequency comb within the range of $3\omega_{\mathrm{m}}\leq\Delta_{\mathrm{p}}\leq 3\omega_{\mathrm{m}}$. Notably, frequency comb can also emerge in the red sideband region of the cavity-2 ($\Delta_{\mathrm{p}}<0$), which is completely absent in Kerr optomechanical combs (for $\kappa<\Omega$[53, 54, 55, 56, 64]. In theory, the pump power could be much lower, but it is constrained by the limitations of the detection accuracy in our experimental setup. Fig. 3c shows an example of the frequency comb at a pump detuning of $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$, with a repetition frequency set by $\omega_{\mathrm{m}}$. As the pump power increases, the number of detectable frequency comb teeth increases accordingly. Meanwhile, the entire set of comb teeth remains symmetrically distributed around $\omega_{\mathrm{c2}}$.
@@
For comparison, we also independently characterized the emergence and evolution of the Kerr optomechanical frequency comb in cavity-2. As shown by the red data points in Fig. 3b, the minimum pump power $P_{\mathrm{p-Kerr}}=11\ \mathrm{dBm}$ occurs at the blue sideband $\Delta_{\mathrm{p}}\approx\omega_{\mathrm{m}}$ because of maximized antidamping, triggering the optomechanical instability most efficiently.
@@
Focusing on the Floquet mode of $m=1$ in Fig. 3b, the minimum generated power for the Floquet frequency comb is approximately reduced by six orders of magnitude compared to Kerr optomechanical combs. Moreover, for the $m=1$ mode, the width of the local minimum is broader and shifts toward $m=0$. This occurs because, at this point, the pump not only serves as the probe of the Floquet frequency comb, but also drives the mechanical oscillator, acting similarly to the mechanism in Kerr optomechanical combs. The interaction between the pump signal and the mechanical oscillator results in a broader local minimum and the shift of this mode.
Effect of modulation source and coherence
@@
Figure 4: Parameter dependence and phase coherence. a Effect of drive and pump power on Floquet cavity frequency comb generation at m=1,0,+1m=-1,0,+1 (Red, purple, blue). The colored regions indicate that the Floquet comb is generated, whereas the white areas represent the absence of a detectable frequency comb. The green dashed line represents the power range selected for subsequent analysis. b Example frequency comb under $\Delta_{\text{p}}=-\omega_{\text{m}}$ in the frequency domain. c Effect of modulation strength on the relative magnitudes of the frequency comb teeth, with error bars representing standard deviations. The gradient-colored points represent the peak values of the corresponding $m$-th comb tooth. d Phasor distribution and temporal evolution of the phase deference $\Delta\phi_{-1,1}(t)$. The amplitude is normalized to its mean value and all carrier frequencies are removed. The color scale in the phasor plot represents the counts per unit area, while the red dashed line in the temporal trace indicates the mean value of $\Delta\phi_{-1,1}(t)$. e Representative time slices of the 200 ms in-phase (II) and quadrature ($Q$) components: an early-time segment from 10 µs to 30 µs after the onset of the recorded trace (blue) and a late-time segment from 199.97 ms to 199.99 ms of the trace (red). The insets provide enlarged views of each window.
@@
We investigate the impact of modulation strength on the Floquet frequency comb by setting the pump tone at $\Delta_{\mathrm{p}}/\omega_{\mathrm{m}}=-1,0,+1$ to match the $m$-th Floquet mode. As shown in Fig. 4a, while varying both the drive power and pump power, for $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$, the frequency comb is generated only when the Floquet modulation is activated, $P_{\mathrm{d}}>P_{\mathrm{dc}}$. The fluctuations in the minimum power value are attributed to measurement uncertainties. Notably, for $\Delta_{\mathrm{p}}=0$, when the pump power is sufficiently high, the frequency comb is generated before $P_{\mathrm{dc}}$. This is the result of the cooperative action of the optomechanical frequency comb under the external driving of two cavities. The results for $\Delta_{\mathrm{p}}=\omega_{\mathrm{m}}$ exhibit similar behavior. At sufficiently high pump power, cavity-2 can independently generate a frequency comb via Kerr optomechanical interactions. Thus, we observe a gradual onset rather than a sharp transition.
It should be noted that, despite the absence of a threshold with respect to the pump tone, a finite total power consumption is still required to sustain self-induced oscillations. Due to the presence of attenuators (approximately 65 dB) between the microwave sources at room temperature and the cavity input port at 10 mK (see Supplementary Note III.B), a typical source power of 10 dBm corresponds to approximately 3.2 nW at the cavity. If we consider the mechanical oscillator itself, the power required to resonantly drive it to a vibrational amplitude of 1 nm is only 93 fW, suggesting that the overall power consumption can be further reduced. To minimize interference from the Kerr optomechanical frequency comb in our analysis, we focus on the pump at $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$. Fig. 4b shows an example of Floquet frequency comb with $P_{\mathrm{d}}=11~\mathrm{dBm}$ and $P_{\mathrm{p}}=0~\mathrm{dBm}$.
@@
We demonstrate the effect of modulation strength for the $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$ case with $P_{\mathrm{p}}=0~\mathrm{dBm}$, shown in Fig. 4c. As the modulation strength increases, the Floquet cavity’s ability to scatter photons into higher-order sidebands strengthens, resulting in a decrease in the intensity of the peak at $m=0$ and an increase in the intensities of peaks at $m=\pm 1,\pm 2,\dots$. With further increase in modulation strength, the intensity of the lower-order sidebands stabilizes and begins to decrease, while the effect of higher-order sidebands becomes more pronounced. At higher $P_{\mathrm{d}}$ values, the intensity of the m=1m=-1 mode exhibits an anomalous increase due to its spectral overlap with the pump tone. We also observe an asymmetry between the two sets of sidebands, which is an intrinsic feature of the Floquet cavity frequency comb resulting from the finite pump detuning ($\Delta_{\mathrm{p}}\neq 0$).
We perform a preliminary verification of the coherence by measuring the phase difference $\Delta\phi_{-1,1}(t)$ obtained from the beat note between the m=1m=-1 and $m=+1$ comb teeth. Fig. 4d shows both the phasor distributions and the temporal evolution of $\Delta\phi_{-1,1}(t)$ under $\omega_{\mathrm{p}}=\omega_{\mathrm{c2}}$. The phasors are tightly clustered and remain confined within a narrow angular sector, rather than being uniformly distributed over the unit circle, indicating that the phase difference between the comb lines is stable. The temporal evolution further demonstrates bounded phase fluctuations without phase diffusion over the full $2\pi$ range. Meanwhile, we recorded the temporal phase evolution of the $m=+1$ and $m=+2$ comb teeth and compared the segments from 10 µs to 30 µs after the start and from 30 µs to 10 µs before the end of the trace, as shown in Fig. 4e. The phase traces at different times overlap, indicating the phase coherence and stability of the comb teeth. Additional verification of coherence was obtained through measurements of phase noise and cross-correlation between individual comb teeth (Supplementary Note V).
@@ Discussion
We have demonstrated that cavity Floquet engineering is a viable option for the generation of a frequency comb. While our current demonstration shows approximately ten comb teeth, this number is not an intrinsic physical limitation of the Floquet mechanism. In the experiment, the drive power was deliberately limited to avoid saturation of the cryogenic and room-temperature microwave amplifiers, which would otherwise introduce distortions and restrict the measurable comb bandwidth. The straightforward approach to increasing the number of comb teeth is to enhance the modulation strength. In our system, this can be achieved either by increasing the single-photon coupling strength to the Floquet cavity ($g_{0,2}$) or by strengthening the external drive applied to the oscillator. Apart from optomechanical driving, alternative excitation methods include capacitive excitation [65, 66], piezoelectric excitation [67], and dielectric force excitation [68], which can further increase the amplitude of the oscillator.
The Floquet cavity frequency comb demonstrated here exhibits a repetition frequency of 9.1 MHz, which matches $\omega_{\text{m}}$. Integrating multiple mechanical oscillators on the same chip or utilizing higher-order vibration modes of the mechanical oscillators can provide additional degrees of freedom for tuning the repetition frequency of the comb. More generally, through appropriate mechanical design, the vibration frequency of mechanical oscillators can span a wide range from the kilohertz to the gigahertz regime [50]. Moreover, techniques such as injection locking [49], feedback control [69, 70], and the use of higher-$Q$ mechanical oscillators [71] provide promising routes further to enhance the stability of the Floquet cavity frequency comb.
Floquet cavity engineering paves the way for the efficient generation and control of frequency combs, rendering them well suited for low-power operation and compatibility with cryogenic quantum systems. For example, in superconducting quantum computing devices, components such as superconducting resonators and qubits cannot operate under high photon occupations[72]. Our system offers an on-chip integrated solution for low-temperature frequency conversion or frequency-division multiplexing [73, 74]. The achieved several-megahertz repetition rate is sufficient to resolve qubits, which typically exhibit kilohertz-scale linewidths [75]. Moreover, Floquet cavity frequency combs can be realized and applied in various systems exhibiting periodic modulation, such as Josephson junction [39, 76], trapped ions [29], quantum dot [77], and molecular or atomic systems [78, 79], particularly in the optical domain. This holds promise for extending the capabilities of frequency combs in high-dimensional quantum states [80], entanglement generation [36, 37], and integrated multi-frequency signal sources across more systems [76, 29].
@@ Methods
here $\psi_{n}(t)=\langle n|\psi,t\rangle$. Combined with Equation (3), the steady-state solution can be expressed as
@@
$\left|u_{n},t\right\rangle$ $=e^{-in\frac{A}{\Omega}\sin\Omega t}|n\rangle.$
We define $\left|u_{nm},t\right\rangle=\left|u_{n},t\right\rangle e^{im\Omega t}=u_{nm}(t)|n\rangle$, with $u_{nm}(t)=e^{-in\frac{A}{\Omega}\sin\Omega t}e^{im\Omega t}$, which satisfies
@@
At this point, we have identified Brillouin regions over the frequency domain, where the single Brillouin region has a width of $\Omega$, and the different Brillouin regions are represented by the index $m$, called the Floquet index. And where $H_{F}=H(t)-i\partial_{t}$ is called the Floquet Hamiltonian, $\left|u_{nm},t\right\rangle$ is the time-periodic Floquet state, and quasi-energy $\varepsilon_{nm}=n\omega_{0}+m\Omega$. The $\left|u_{n},t\right\rangle$ can also be written as $\left|u_{n0},t\right\rangle$. Under the complete basis vector set $\left.\left\{\left|n_{mn}\right\rangle\right\rangle\right\}$ with $\left|n_{m}\right\rangle=|n\rangle e^{im\Omega t}$, the matrix elements of the Floquet Hamiltonian are
@@
where $H^{[0]}=\sum_{n}n\omega_{c}|n\rangle\langle n|$, $H^{[\pm 1]}=\sum_{n}n\frac{A}{2}|n\rangle\langle n|$, and $H^{[l]}=0$ for $|l|>1$. Thus, the Equation (7) can be further expressed as
@@
$H_{F}^{(n)}=\left[\begin{array}[]{ccccc}\ddots&\vdots&\vdots&\vdots&\iddots\\ \cdots&n\omega_{0}-\Omega&n\frac{A}{2}&0&\cdots\\ \cdots&n\frac{A}{2}&n\omega_{0}&n\frac{A}{2}&\cdots\\ \cdots&0&n\frac{A}{2}&n\omega_{0}+\Omega&\cdots\\ \iddots&\vdots&\vdots&\vdots&\ddots\end{array}\right]\text{, }$ (10)
which satisfies $H_{F}^{(n)}\left|u_{nm}\right\rangle\rangle=\varepsilon_{nm}\left|u_{nm}\right\rangle\rangle$ according to the Equation (6). It can be observed that each Fock state energy level splits into $2m+1$ sublevels, consistent with the periodic nature of the Floquet Hamiltonian. The off-diagonal elements, proportional to $A/2$, characterize the coupling strength between adjacent Floquet states, thereby determining the transition rates between different energy levels. This coupling reflects the influence of the periodic driving field and results in transitions mediated by the exchange of discrete energy quanta of size $\Omega$. For more detailed derivations, please refer to the Supplementary Note I.B.
@@
In the blue-detuned regime, the optomechanical damping rate $\gamma_{\mathrm{opt}}$ is negative. Initially, this results in an increase in the oscillator’s effective temperature. As the overall damping rate $\gamma_{\mathrm{m}}+\gamma_{\mathrm{opt}}$ becomes negative, any small initial fluctuation will grow exponentially over time, eventually reaching a steady-state regime. These are referred to as self-induced optomechanical oscillations, where the mechanical oscillation amplitude stabilizes at a fixed value $X_{\mathrm{m}}$. The governing dynamical equations are given by
@@
Ignoring in the initial phase $\phi_{\mathrm{m}}$, we have $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t)+X_{\mathrm{0}}$. In the steady state, we define an amplitude-dependent effective optomechanical damping rate as
@@
$\left\{\begin{aligned} m_{\mathrm{eff}}\omega^{2}_{\mathrm{m}}X_{\mathrm{0}}&=\left\langle F\right\rangle,\\ \gamma_{\mathrm{m}}+\gamma_{\mathrm{opt}}&=0.\end{aligned}\right$. (14)
@@
and the global phase $\varphi(t)=(g_{\mathrm{01}}X_{\mathrm{m}}/x_{\mathrm{zpf}}\omega_{\mathrm{m}})\sin(\omega_{\mathrm{m}}t)$. By substituting this Fourier series into Equation (14), we obtain the implicit equation for $X_{\mathrm{0}}$ and $X_{\mathrm{m}}$:
$\left\{\begin{aligned} X_{\mathrm{0}}&=\frac{\hbar g_{\mathrm{01}}}{x_{\mathrm{zpf}}m_{\mathrm{eff}}\omega^{2}_{\mathrm{m}}}\sum_{k}|a_{k}|^{2},\\ X_{\mathrm{m}}&=\frac{2\hbar g_{\mathrm{01}}}{x_{\mathrm{zpf}}m_{\mathrm{eff}}\omega_{\mathrm{m}}\gamma_{\mathrm{m}}}\mathbf{Im}{\sum_{k}a_{k}^{*}a_{k+1}}.\end{aligned}\right$. (16)
@@
The data supporting the findings of this study are available within the article and its Supplementary Information files. Data are available from the corresponding authors upon request. Source data are provided with this paper.